Droperidol is often used to treat postoperative nausea. an order to administrate 0.275 mg of Droperidol by IV over 3.0 hours. The IV bag contains 2.5 mg/mL Droperidol. What should you set the flow rate to be in milliliters per hour? How many mg of Droperidol will be administered after 24 hours?
Answer #1

Ans) Order:  to administrate 0.275 mg of Droperidol by IV over 3.0 hours.

Available: The IV bag contains 2.5 mg/mL Droperidol.

2.5 mg = 1 ml

0.275 mg = x

x = 0.275/2.5 = 0.11 ml

Flow rate: Total volume/time in hr

ml/hr = 0.11/3 = 0.036 ml/hr

0.036× 24 = 0.864 ml/hr in 24 hours.